### v1

@@ 4.2 Main Results
ZMT demonstrated the most outstanding performance in the qualification examination for TCM practitioners.
As shown in Table 4, it achieved the highest scores in 10 out of 11 test subsets. With the inclusion of updated test items (such as the 2024 official exam questions and Human-Crafted items), ZMT’s advantage has further expanded. Because these items carry a lower risk of data leakage, they provide a more authentic reflection of the model’s true capabilities, thereby underscoring ZMT’s robustness and authority in the field of TCM.
@@
Although the untrained Deepseek-R1 already achieved an Overall score of 89.04, it still lagged behind the state-of-the-art models. After training, however, ZMT’s score increased to 96.17, representing a gain of 7.13 points. This indicates that our training corpus—comprising textbooks and mock exam questions—provides comprehensive coverage of knowledge points in TCM and delivers targeted enhancement of the model’s examination performance.
@@ 5 Analysis
In this section, we conduct a comprehensive analysis to answer the three research questions RQ1: How does the model perform across different sub-domains or tasks within TCM? (Sec 5.1) RQ2: Is the SI-CoTE training approach universally effective across different model architectures? (Sec 5.2) RQ3: Through case studies, does the model genuinely acquire knowledge specific to TCM? (Sec 5.3)
@@ 5.1 Analysis of Model Performance Across Various Aspects of TCM
Figure 3: Performance distribution of ZMT and baseline models across four evaluation units. ZMT exhibits the most balanced performance profile with stable scores on all units. In contrast, baseline models demonstrate significant performance variations, particularly weaker on Units 3 and 4 compared to Unit 2.
As shown in Table 2, TCM-Eval is divided into four units, each designed to assess a distinct capability within the domain of TCM. We have separately evaluated and compared the performance of various models across these four units. Figure 3 reveals a significant variance in the models’ abilities across different dimensions of TCM. Specifically, the scores in Unit 4 are substantially lower than those in Unit 2, indicating that existing models have a notable deficiency in clinical application skills. In contrast, our proposed ZMT model not only surpasses other models in every unit but also demonstrates the most balanced proficiency, achieving the lowest score variance. This superior and well-rounded performance is attributed to the comprehensive coverage of knowledge points during our model’s training phase.
@@ 5.2 Generalizability of the SI-CoTE
To ascertain the efficacy of the CoT augmented training data from Section 3.2, we conducted fine-tuning experiments on open-source models with known deficiencies in the TCM domain. As presented in Table 5, fine-tuning the Qwen3 and Deepseek-Distill series with our data yielded significant performance enhancements. The improvements were particularly pronounced for smaller models, which saw performance boosts of up to 135%, rivaling the capabilities of API-based models. These findings underscore the high quality and broad coverage of the data produced by our SI-CoTE iterative method, confirming its effectiveness in advancing model performance in the TCM field.
@@ 5.3 Case Study
This section presents a qualitative, side-by-side comparison of DeepSeek-R1 and ZMT (Figure 4), highlighting key differences in their approach to and understanding of TCM. The question shown in the figure combines textbook knowledge with clinical diagnostic experience. It is evident that ZMT has a broader knowledge base within the scope of TCM compared to DeepSeek-R1. ZMT demonstrates a better grasp of textbook knowledge and integrates real clinical experience. Additionally, ZMT provides more concise and accurate answers, reflecting its greater expertise in the field of TCM.
@@ 6 Conclusion
This work addresses critical gaps in LLM applications for TCM. We introduce TCM-Eval, the first dynamic, expert-validated benchmark sourced from national licensing examinations, establishing a rigorous ’gold standard’ for robust LLM evaluation. To foster advanced reasoning, our novel Self-Iterative Chain-of-Thought Enhancement (SI-CoTE) framework enables autonomous CoT generation and validation, driving a virtuous cycle of data and model co-evolution. Leveraging this, our SOTA LLM, ZMT, achieves an unprecedented 96.32% on TCM-Eval, far surpassing human practitioner pass rates and setting a new benchmark for expert-level knowledge and reasoning. By open-sourcing these resources and maintaining a dynamic public leaderboard, we aim to catalyze future research and accelerate TCM’s AI-driven modernization.

### v2

@@ 4.2 Main Results
ZMT-M1 demonstrated the most outstanding performance in the qualification examination for TCM practitioners.
As shown in Table 4, it achieved the highest scores in 10 out of 11 test subsets. With the inclusion of updated test items (such as the 2024 official exam questions and Human-Crafted items), ZMT-M1’s advantage has further expanded. Because these items carry a lower risk of data leakage, they provide a more authentic reflection of the model’s true capabilities, thereby underscoring ZMT-M1’s robustness and authority in the field of TCM.
@@
Although the untrained Deepseek-R1 already achieved an Overall score of 89.04, it still lagged behind the state-of-the-art models. After training, however, ZMT-M1’s score increased to 96.17, representing a gain of 7.13 points. This indicates that our training corpus—comprising textbooks and mock exam questions—provides comprehensive coverage of knowledge points in Traditional Chinese Medicine and delivers targeted enhancement of the model’s examination performance.
@@ 5 Analysis
In this section, we conduct a comprehensive analysis to answer the three research questions RQ1: How does the model perform across different sub-domains or tasks within traditional Chinese medicine? (Sec 5.1) RQ2: Is the SI-CoTE training approach universally effective across different model architectures? (Sec 5.2) RQ3: Through case studies, does the model genuinely acquire knowledge specific to traditional Chinese medicine after training? (Sec 5.3)
@@ 5.1 Analysis of Model Performance Across Various Aspects of TCM
Figure 3: Performance distribution of ZMT-M1 and baseline models across four evaluation units. ZMT-M1 exhibits the most balanced performance profile with stable scores on all units. In contrast, baseline models demonstrate significant performance variations, particularly weaker on Units 3 and 4 compared to Unit 2.
As shown in Table 2, TCM-Eval is divided into four units, each designed to assess a distinct capability within the domain of TCM. We have separately evaluated and compared the performance of various models across these four units. Figure 3 reveals a significant variance in the models’ abilities across different dimensions of TCM. Specifically, the scores in Unit 4 are substantially lower than those in Unit 2, indicating that existing models have a notable deficiency in clinical application skills. In contrast, our proposed ZMT-M1 model not only surpasses other models in every unit but also demonstrates the most balanced proficiency, achieving the lowest score variance. This superior and well-rounded performance is attributed to the comprehensive coverage of knowledge points during our model’s training phase.
@@ 5.2 Generalizability of the SI-CoTE
To ascertain the efficacy of the CoT augmented training data from Section LABEL:sec:method, we conducted fine-tuning experiments on open-source models with known deficiencies in the TCM domain. As presented in Table 5, fine-tuning the Qwen3 and Deepseek-Distill series with our data yielded significant performance enhancements. The improvements were particularly pronounced for smaller models, which saw performance boosts of up to 135%, rivaling the capabilities of API-based models. These findings underscore the high quality and broad coverage of the data produced by our SI-CoTE iterative method, confirming its effectiveness in advancing model performance in the TCM field.
@@ 5.3 Case Study
This section presents a qualitative, side-by-side comparison of DeepSeek-R1 and ZMT-M1 (Figure 4), highlighting key differences in their approach to and understanding of TCM. The question shown in the figure combines textbook knowledge with clinical diagnostic experience. It is evident that ZMT-M1 has a broader knowledge base within the scope of TCM compared to DeepSeek-R1. ZMT-M1 demonstrates a better grasp of textbook knowledge and integrates real clinical experience. Additionally, ZMT-M1 provides more concise and accurate answers, reflecting its greater expertise in the field of TCM.
@@ 6 Conclusion
This work addresses critical gaps in LLM applications for TCM. We introduce TCM-Eval, the first dynamic, expert-validated benchmark sourced from national licensing examinations, establishing a rigorous ’gold standard’ for robust LLM evaluation. To foster advanced reasoning, our novel Self-Iterative Chain-of-Thought Enhancement (SI-CoTE) framework enables autonomous CoT generation and validation, driving a virtuous cycle of data and model co-evolution. Leveraging this, our SOTA LLM, ZMT-M1, achieves an unprecedented 96.32% on TCM-Eval, far surpassing human practitioner pass rates and setting a new benchmark for expert-level knowledge and reasoning. By open-sourcing these resources and maintaining a dynamic public leaderboard, we aim to catalyze future research and accelerate TCM’s AI-driven modernization.